# Autonomous Constellation Fault Monitoring with Inter-satellite Links: A Rigidity-Based Approach

### v1

@@ 4.3 Case 2: Elliptical Lunar Frozen Orbit Constellation
(b) Number of 6-cliques for different time epochs. The satellite near the perilune has few or zero 6-cliques containing itself due to the occultation by the Moon.
Figure 9: Case 2: ELFO constellation around Moon with 12 satellites.
Figure 10: Case 2 (ELFO constellation): The detection results for different fault magnitude and the detection length (DL). The number of faults is fixed to one. The detector using the fixed threshold is shown in a dashed line, while the detector using the trained neural network is shown in a solid line.
(a) Detection results with fixed threshold
(b) Detection results with neural network predictor
Figure 11: Case 2 (ELFO constellation): Detection results (z-axis) shown as a 3D plot for different fault magnitudes (x-axis), number of detection timesteps (y-axis), and number of faults (planes).
@@ 5 Conclusion
In this paper, we introduced a fault detection framework for autonomous satellite constellations using inter-satellite range (ISR) measurements. Our fault detection approach is based solely on the ISR measurements, and does not require the availability of precise ephemeris information. Compared to traditional fault monitoring methods that require ground station monitoring or precise ephemeris, our proposed method is suited for lunar and Martian environments where dedicated monitoring stations on the planetary surface would not be available.
In the paper, we start by providing mathematical proofs of several properties related to the rank of the EDM and GCEDM. Based on these properties, we demonstrated that satellite fault can be detected by analyzing the singular values and the singular vectors of the geometric centered Euclidean Distance Matrix (GCEDM) derived from ISR measurements. By combining a neural network predictor and the clique finding algorithm, the proposed fault detection framework can adapt the detection threshold to dynamic geometries and topologies, enhancing the robustness and reliability of fault detection. Through simulations of constellations around Mars and the Moon, we demonstrated the effectiveness of our fault detection framework in various configurations and organized the relationship between the environmental parameters and the detection metric.
This research marks a step towards establishing reliable operation of satellite constellations in extraterrestrial environments, which is the key architecture to support future exploration missions on the Moons or Mars. Future work will focus on refining the neural network predictor to accommodate more diverse and complex satellite constellations and combining the proposed algorithms with residual-based fault detection methods to further enhance the robustness and scalability of autonomous satellite fault detection.
@@ acknowledgements
We acknowledge Derek Knowles for reviewing the paper. This material is based upon work supported by the National Science Foundation under Grant No. DGE-1656518 and CNS-2006162.
@@
$\mathbf{D}^{n,d,m}_{ij}$ $=(\|\bm{x}_{i}-\bm{x}_{j}\|+b_{ij})^{2}=\|\bm{x}_{i}-\bm{x}_{j}\|^{2}+2\|\bm{x}_{i}-\bm{x}_{j}\|b_{ij}+b_{ij}^{2}$ (16)
$=\|\bm{x}_{i}-\bm{x}_{j}\|^{2}+b_{ij}\left(2\|\bm{x}_{i}-\bm{x}_{j}\|+b_{ij}\right)$
$=\|\bm{x}_{i}-\bm{x}_{j}\|^{2}+b_{ij}\left(s_{ij}+b_{ij}\right)$
Where $s_{ij}$ is two times the elementwise square root of the corresponding entry in the EDM. As a matrix, we can write $\mathbf{S}^{n,d}_{ij}=s_{ij}$. We can write the bias term $\mathbf{B}^{n,m}_{ij}=b_{ij}$ as a matrix as well. Both $\mathbf{S}^{n,d}$ and $\mathbf{B}^{n,m}$ are real, symmetric matrices. Then, we arrive at Equation (17).
$\mathbf{D}^{n,d,m}=\mathbf{D}^{n,d,0}+\mathbf{B}^{n,m}\circ(\mathbf{S}^{n,d}+\mathbf{B}^{n,m})$ (17)
@@
$\mathbf{B}^{n,\{2,3,4\}}=\bar{b}\begin{bmatrix}0&1&1&1&0&\cdots&0\\ 1&0&2&2&1&\cdots&1\\ 1&2&0&2&1&\cdots&1\\ 1&2&2&0&1&\cdots&1\\ 0&1&1&1&0&\cdots&0\\ \vdots&\vdots&\vdots&\vdots&\vdots&\ddots&\vdots\\ 0&1&1&1&0&\cdots&0\end{bmatrix}=\bar{b}\begin{bmatrix}0&1&1&1\\ 1&0&2&2\\ 1&2&0&2\\ 1&2&2&0\\ 0&1&1&1\\ \vdots&\vdots&\vdots&\vdots\\ 0&1&1&1\\ \end{bmatrix}\begin{bmatrix}1&0&0&0&1&\cdots&1\\ 0&1&0&0&0&\cdots&0\\ 0&0&1&0&0&\cdots&0\\ 0&0&0&1&0&\cdots&0\\ \end{bmatrix}$ (18)
@@
In the rank-decomposition of $\mathbf{B}^{n,m}$, the column matrix has the first column has a one in each fault entry and a zero otherwise. This column spans all the $n-m$ fault-free columns. The remaining $m$ columns correspond to the columns in $\mathbf{B}^{n,m}$ with faults and have a zero at the fault entry. $\mathbf{B}^{n,m}$ achieves full rank when the number of faults reaches $m=n-1$. In general, $\mathrm{rank}(\mathbf{B}^{n,m})=\min(1+m,n)$ if $m>0$ and $\mathrm{rank}(\mathbf{B}^{n,0})=0$. By similar argument, $\mathrm{rank}(\mathbf{B}^{n,m}\circ\mathbf{B}^{n,m})=\mathrm{rank}(\mathbf{B}^{n,m})$ since we will have one column indicating the fault entries and $m$ columns associated with the columns of $\mathbf{B}^{n,m}\circ\mathbf{B}^{n,m}$ with faults, which are the same columns as $\mathbf{B}^{n,m}$. This overall rank relationship often holds even if the faults are of different non-zero magnitudes. However, if the faults cancel out, the rank will drop. For example, in the case above, if $b_{2}=2\bar{b}$ and $b_{3}=b_{4}=-\bar{b}$, then $\mathrm{rank}(\mathbf{B}^{n,m})=3$. Therefore, $\mathrm{rank}(\mathbf{B}^{n,m})\leq\min(1+m,n)$ for arbitrary fault sizes. Nevertheless, the term $\mathrm{rank}(\mathbf{B}^{n,m}\circ\mathbf{B}^{n,m})$ can still be rank $m+1$ even if $\mathrm{rank}(\mathbf{B}^{n,m})<m+1$ since the squaring can remove the linear cancellation.
@@
However, the span of the columns of $\mathbf{B}^{n,m}\circ\mathbf{S}^{n,d}$ will generally span the columns $\mathbf{B}^{n,m}\circ\mathbf{B}^{n,m}$. Again, for the sake of illustration, we extend the example above with faults at satellites $k=\{2,3,4\}$ for $n>7$, without loss of generality.
$\mathbf{B}^{n,\{2,3,4\}}\circ\mathbf{S}^{n,d}$ $=\bar{b}\begin{bmatrix}0&s_{12}&s_{13}&s_{14}&0&\cdots&0\\ s_{12}&0&2s_{23}&2s_{24}&s_{25}&\cdots&s_{2n}\\ s_{13}&2s_{23}&0&2s_{34}&s_{35}&\cdots&s_{3n}\\ s_{14}&2s_{24}&2s_{34}&0&s_{45}&\cdots&s_{4n}\\ 0&s_{25}&s_{35}&s_{45}&0&\cdots&0\\ \vdots&\vdots&\vdots&\vdots&\vdots&\ddots&\vdots\\ 0&s_{2n}&s_{3n}&s_{4n}&0&\cdots&0\end{bmatrix}$ (19)
$=\bar{b}\begin{bmatrix}0&s_{12}&s_{13}&s_{14}&0&0\\ s_{12}&0&2s_{23}&2s_{24}&s_{25}&s_{26}\\ s_{13}&2s_{23}&0&2s_{34}&s_{35}&s_{36}\\ s_{14}&2s_{24}&2s_{34}&0&s_{45}&s_{46}\\ 0&s_{25}&s_{35}&s_{45}&0&0\\ \vdots&\vdots&\vdots&\vdots&\vdots&\vdots\\ 0&s_{2n}&s_{3n}&s_{4n}&0&0\\ \end{bmatrix}\begin{bmatrix}1&0&0&0&0&0&\Delta_{17}&\cdots&\Delta_{1n}\\ 0&1&0&0&0&0&0&\cdots&0\\ 0&0&1&0&0&0&0&\cdots&0\\ 0&0&0&1&0&0&0&\cdots&0\\ 0&0&0&0&1&0&\Delta_{57}&\cdots&\Delta_{5n}\\ 0&0&0&0&0&1&\Delta_{67}&\cdots&\Delta_{6n}\\ \end{bmatrix}$
@@
The columns needed to span $\mathbf{B}^{n,m}\circ\mathbf{S}^{n,d}$ are the same as those needed for $\mathbf{B}^{n,m}\circ\mathbf{B}^{n,m}$, with the same sparsity pattern. Therefore, we will have
$\mathrm{rank}(\mathbf{B}^{n,m}\circ(\mathbf{S}^{n,d}+\mathbf{B}^{n,m}))=\mathrm{rank}(\mathbf{B}^{n,m}\circ\mathbf{S}^{n,d})\leq\min(2m,n)$ (20)
@@
$=-\frac{1}{2}J^{n}(\mathbf{D}^{n,d,0}+\mathbf{B}^{n,m}\circ(\mathbf{S}^{n,d}+\mathbf{B}^{n,m}))J^{n}$
$=-\frac{1}{2}J^{n}(\mathbf{1}\,\mathrm{diag}(\mathbf{X}^{\top}\mathbf{X})^{\top}-2\mathbf{X}^{\top}\mathbf{X}+\mathrm{diag}(\mathbf{X}^{\top}\mathbf{X})\mathbf{1}^{\top})J^{n}-\frac{1}{2}J^{n}(\mathbf{B}^{n,m}\circ(\mathbf{S}^{n,d}+\mathbf{B}^{n,m}))J^{n}$
@@
The remaining term is $\frac{1}{2}J^{n}(\mathbf{B}^{n,m}\circ(\mathbf{S}^{n,d}+\mathbf{B}^{n,m}))J^{n}$. By rank inequality,
$\mathrm{rank}(J^{n}(\mathbf{B}^{n,m}\circ(\mathbf{S}^{n,d}+\mathbf{B}^{n,m}))J^{n})\leq\min(\mathrm{rank}(\mathbf{B}^{n,m}\circ(\mathbf{S}^{n,d}+\mathbf{B}^{n,m})),\mathrm{rank}(J^{n}))=\min(2m,n-1)$ (26)
@@
Appendix C Proof of Proposition 2.2.1
@@
In terms of rank, the noise acts as many small faults on each satellite, with the same sparsity structure as $\mathbf{B}^{n,m}$ in Proposition 2.1, almost surely with $m=n$ since there is zero probability mass that the randomly sampled noise is exactly zero or exactly cancel out. Using Proposition 2.2,

### v2

@@ 5.2 Simulation Result
Figure 10 and Figure 11 show the relationship between the test statistic threshold, detection length, and the three performance metrics. In general, TPR improves as the test statistic threshold $\bar{\gamma}_{\text{test}}$ decreases and detection length $DL$ increases. On the contrary, FPR improves as the threshold increases and detection length decreases. This is because lowering the threshold and increasing the detection length both increase the number of subgraphs detected as faults, which increases the probability of detecting a faulty satellite but also increases the probability of detecting normal satellites as faults. The P4 value, which shows the balance, is maximized at higher thresholds and longer DL when fault magnitudes are smaller, and vice versa for larger fault magnitudes.
Figure 12 shows the relationship between the number of faults and the three performance metrics, for a fixed detection threshold and detection length. In general, TPR, FPR, and $P_{4}$ value all improve when the number of faults is smaller, except for TPR and $P_{4}$ in small fault magnitude cases where it becomes easier to detect at least one fault out of multiple faults. Our proposed algorithms removes fault satellites in a greedy manner, so when there are multiple fault satellites, the larger perturbation in the 4th and 5th singular value and the 4th singular vector makes it difficult to determine the correct set of fault satellites. We could potentially improve our detection algorithm by also looking at the singular values of larger indices (e.g., 6th and 7th), but that would require using larger subgraphs for detection.
Table 2: Fault detection results for different numbers of faults, fault magnitude, fault detection threshold, and detection length (DL). The best TPR, FPR, and P4 values for each (number of faults, fault magnitude) pair are highlighted in blue. All three metrics improve when the fault magnitude is larger, and the number of faults is smaller. TPR improves as the threshold decreases and detection length increases, and FPR improves as the threshold increases and detection length decreases. The $P_{4}$ values are better at higher thresholds and longer DL at smaller fault magnitudes, and vice versa for larger fault magnitudes.
    Faults Threshold DL     True Positive Rate     False Positive Rate     P4 value
  $\bar{\gamma}_{\text{test}}$     Fault Magnitude [m]     Fault Magnitude [m]     Fault Magnitude [m]
[…]
    1 99.9 1     0.006 0.374 0.864     0.001 0.000 0.001     0.023 0.698 0.955
  2     0.042 0.582 0.896     0.001 0.002 0.002     0.147 0.833 0.964
  3     0.048 0.742 0.908     0.002 0.002 0.001     0.165 0.908 0.969
  5     0.124 0.854 0.918     0.005 0.001 0.003     0.345 0.951 0.968
  99 1     0.146 0.776 0.900     0.015 0.005 0.006     0.362 0.911 0.952
  2     0.450 0.866 0.914     0.023 0.014 0.013     0.682 0.917 0.936
  3     0.530 0.874 0.924     0.032 0.026 0.019     0.712 0.888 0.922
  5     0.744 0.892 0.926     0.050 0.041 0.039     0.774 0.854 0.871
  95 1     0.478 0.866 0.908     0.055 0.039 0.037     0.618 0.852 0.870
  2     0.712 0.896 0.922     0.076 0.069 0.075     0.703 0.792 0.790
  3     0.758 0.888 0.918     0.098 0.096 0.098     0.681 0.736 0.743
  5     0.808 0.906 0.922     0.108 0.115 0.116     0.685 0.709 0.713
    2 99.9 1     0.109 0.328 0.813     0.004 0.025 0.005     0.319 0.606 0.929
  2     0.173 0.458 0.853     0.007 0.026 0.005     0.436 0.714 0.944
  3     0.216 0.515 0.863     0.009 0.027 0.007     0.500 0.750 0.945
  5     0.250 0.608 0.884     0.018 0.033 0.005     0.531 0.795 0.956
  99 1     0.246 0.603 0.854     0.022 0.022 0.009     0.521 0.810 0.939
  2     0.376 0.721 0.875     0.030 0.028 0.010     0.643 0.857 0.944
  3     0.439 0.774 0.880     0.045 0.027 0.015     0.672 0.881 0.938
  5     0.537 0.802 0.896     0.060 0.038 0.020     0.716 0.877 0.937
  95 1     0.417 0.775 0.865     0.055 0.032 0.023     0.641 0.874 0.921
  2     0.581 0.815 0.875     0.071 0.056 0.045     0.727 0.856 0.894
  3     0.627 0.841 0.885     0.090 0.067 0.066     0.729 0.852 0.870
  5     0.664 0.850 0.886     0.106 0.093 0.090     0.728 0.823 0.840
    3 99.9 1     0.161 0.277 0.620     0.010 0.050 0.043     0.414 0.537 0.800
  2     0.197 0.338 0.659     0.018 0.060 0.046     0.463 0.589 0.817
  3     0.214 0.379 0.685     0.024 0.063 0.046     0.482 0.622 0.830
  5     0.257 0.410 0.713     0.030 0.078 0.048     0.531 0.634 0.841
  99 1     0.233 0.428 0.675     0.032 0.055 0.047     0.501 0.669 0.824
  2     0.297 0.513 0.721     0.054 0.066 0.045     0.554 0.717 0.848
  3     0.330 0.566 0.732     0.068 0.068 0.049     0.574 0.747 0.849
  5     0.399 0.597 0.739     0.078 0.075 0.052     0.626 0.758 0.849
  95 1     0.337 0.572 0.689     0.071 0.067 0.050     0.579 0.751 0.828
  2     0.409 0.625 0.725     0.101 0.074 0.052     0.614 0.774 0.843
  3     0.459 0.672 0.734     0.107 0.074 0.061     0.645 0.797 0.838
  5     0.497 0.671 0.729     0.112 0.089 0.070     0.666 0.783 0.828
Figure 10: The detection results for different fault magnitudes and the test statistic threshold $\bar{\gamma}_{\text{test}}$. The number of faults and the detection length is fixed to 1. Lowering the test statistic threshold results in larger TPR and lower FPR. The $P_{4}$ score is larger at smaller thresholds for smaller fault magnitudes, and at larger thresholds for larger fault magnitudes.
Figure 11: The detection results for different fault magnitudes and the detection length $DL$. The test statistic threshold is fixed to $\bar{\gamma}_{\text{test}}^{99}$, and the number of faults is fixed to 1. Increasing the $DL$ results in a larger TPR and lower FPR. The $P_{4}$ score is larger at longer $DL$ for smaller fault magnitudes, and at shorter $DL$ for larger fault magnitudes.
Figure 12: The detection results for different fault magnitudes and number of fault satellites. The test statistic threshold is fixed to $\bar{\gamma}_{\text{test}}^{99}$, and the detection length is fixed to 1. In general, increasing the number of faults results in lower TPR and higher FPR.
@@ 6 Conclusion
In this paper, we introduced a fault detection framework for autonomous satellite constellations using inter-satellite range (ISR) measurements. Our fault detection approach is based solely on the ISR measurements, and does not require the availability of precise ephemeris information. Compared to traditional fault monitoring methods that require ground station monitoring or precise ephemeris, our proposed method is suited for lunar environments where dedicated monitoring stations on the planetary surface would not be available.
In the paper, we start by providing mathematical proofs of the sufficient and necessary conditions of the graph topology required for fault detection. Next, we introduce several properties of the rank of the EDM and GCEDM. Based on these properties, we demonstrated that satellite fault can be detected by analyzing the singular values and the singular vectors of the geometric centered Euclidean Distance Matrix (GCEDM) derived from ISR measurements, using vertex-redundantly rigid graphs. By using multiple vertex-redundantly rigid subgraphs found by a clique-finding algorithm to detect faults , the proposed fault detection framework can be used for dynamic geometries and graph topologies, enhancing the robustness and reliability of fault detection. Through simulations of an ELFO constellation around the Moon, we demonstrated the effectiveness of our fault detection framework and organized the relationship between the environmental parameters, hyperparameters, and detection performance.
This research marks a step towards establishing reliable operation of satellite constellations in extraterrestrial environments, which is the key architecture to support future exploration missions on the Moon. Future work will focus on developing a model that can predict the tails of the non-fault test statistics distribution from the ranges, selecting effective subgraphs to reduce computation, looking into distributed implementation, and combining the proposed algorithms with residual-based fault detection methods to further enhance the robustness of the algorithm.
@@ acknowledgements
We acknowledge Derek Knowles for reviewing the paper. This material is based upon work supported by The Nakajima Foundation and the National Science Foundation under Grant No. DGE-1656518 and CNS-2006162.
@@
$\mathbf{D}^{n,d,m}_{ij}$ $=(\|\bm{x}_{i}-\bm{x}_{j}\|+f_{ij})^{2}=\|\bm{x}_{i}-\bm{x}_{j}\|^{2}+2\|\bm{x}_{i}-\bm{x}_{j}\|f_{ij}+f_{ij}^{2}$ (14)
$=\|\bm{x}_{i}-\bm{x}_{j}\|^{2}+f_{ij}\left(2\|\bm{x}_{i}-\bm{x}_{j}\|+f_{ij}\right)$
$=\|\bm{x}_{i}-\bm{x}_{j}\|^{2}+f_{ij}\left(s_{ij}+f_{ij}\right)$
Where $s_{ij}$ is two times the elementwise square root of the corresponding entry in the EDM. As a matrix, we can write $\mathbf{S}^{n,d}_{ij}=s_{ij}$. We can write the bias term $\mathbf{F}^{n,m}_{ij}=b_{ij}$ as a matrix as well. Both $\mathbf{S}^{n,d}$ and $\mathbf{F}^{n,m}$ are real, symmetric matrices. Then, we arrive at Equation (15).
$\mathbf{D}^{n,d,m}=\mathbf{D}^{n,d,0}+\mathbf{F}^{n,m}\circ(\mathbf{S}^{n,d}+\mathbf{F}^{n,m})$ (15)
@@
$\mathbf{F}^{n,\{2,3,4\}}=\bar{f}\begin{bmatrix}0&1&1&1&0&\cdots&0\\ 1&0&2&2&1&\cdots&1\\ 1&2&0&2&1&\cdots&1\\ 1&2&2&0&1&\cdots&1\\ 0&1&1&1&0&\cdots&0\\ \vdots&\vdots&\vdots&\vdots&\vdots&\ddots&\vdots\\ 0&1&1&1&0&\cdots&0\end{bmatrix}=\bar{f}\begin{bmatrix}0&1&1&1\\ 1&0&2&2\\ 1&2&0&2\\ 1&2&2&0\\ 0&1&1&1\\ \vdots&\vdots&\vdots&\vdots\\ 0&1&1&1\\ \end{bmatrix}\begin{bmatrix}1&0&0&0&1&\cdots&1\\ 0&1&0&0&0&\cdots&0\\ 0&0&1&0&0&\cdots&0\\ 0&0&0&1&0&\cdots&0\\ \end{bmatrix}$ (16)
@@
In the rank-decomposition of $\mathbf{F}^{n,m}$, the column matrix has the first column has a one in each fault entry and a zero otherwise. This column spans all the $n-m$ fault-free columns. The remaining $m$ columns correspond to the columns in $\mathbf{F}^{n,m}$ with faults and have a zero at the fault entry. $\mathbf{F}^{n,m}$ achieves full rank when the number of faults reaches $m=n-1$. In general, $\mathrm{rank}(\mathbf{F}^{n,m})=\min(1+m,n)$ if $m>0$ and $\mathrm{rank}(\mathbf{F}^{n,0})=0$. By similar argument, $\mathrm{rank}(\mathbf{F}^{n,m}\circ\mathbf{F}^{n,m})=\mathrm{rank}(\mathbf{F}^{n,m})$ since we will have one column indicating the fault entries and $m$ columns associated with the columns of $\mathbf{F}^{n,m}\circ\mathbf{F}^{n,m}$ with faults, which are the same columns as $\mathbf{F}^{n,m}$. This overall rank relationship often holds even if the faults are of different non-zero magnitudes. However, if the faults cancel out, the rank will drop. For example, in the case above, if $b_{2}=2\bar{f}$ and $b_{3}=b_{4}=-\bar{f}$, then $\mathrm{rank}(\mathbf{F}^{n,m})=3$. Therefore, $\mathrm{rank}(\mathbf{F}^{n,m})\leq\min(1+m,n)$ for arbitrary fault sizes. Nevertheless, the term $\mathrm{rank}(\mathbf{F}^{n,m}\circ\mathbf{F}^{n,m})$ can still be rank $m+1$ even if $\mathrm{rank}(\mathbf{F}^{n,m})<m+1$ since the squaring can remove the linear cancellation.
@@
However, the span of the columns of $\mathbf{F}^{n,m}\circ\mathbf{S}^{n,d}$ will generally span the columns $\mathbf{F}^{n,m}\circ\mathbf{F}^{n,m}$. Again, for the sake of illustration, we extend the example above with faults at satellites $k=\{2,3,4\}$ for $n>7$, without loss of generality.
$\mathbf{F}^{n,\{2,3,4\}}\circ\mathbf{S}^{n,d}$ $=\bar{f}\begin{bmatrix}0&s_{12}&s_{13}&s_{14}&0&\cdots&0\\ s_{12}&0&2s_{23}&2s_{24}&s_{25}&\cdots&s_{2n}\\ s_{13}&2s_{23}&0&2s_{34}&s_{35}&\cdots&s_{3n}\\ s_{14}&2s_{24}&2s_{34}&0&s_{45}&\cdots&s_{4n}\\ 0&s_{25}&s_{35}&s_{45}&0&\cdots&0\\ \vdots&\vdots&\vdots&\vdots&\vdots&\ddots&\vdots\\ 0&s_{2n}&s_{3n}&s_{4n}&0&\cdots&0\end{bmatrix}$ (17)
$=\bar{f}\begin{bmatrix}0&s_{12}&s_{13}&s_{14}&0&0\\ s_{12}&0&2s_{23}&2s_{24}&s_{25}&s_{26}\\ s_{13}&2s_{23}&0&2s_{34}&s_{35}&s_{36}\\ s_{14}&2s_{24}&2s_{34}&0&s_{45}&s_{46}\\ 0&s_{25}&s_{35}&s_{45}&0&0\\ \vdots&\vdots&\vdots&\vdots&\vdots&\vdots\\ 0&s_{2n}&s_{3n}&s_{4n}&0&0\\ \end{bmatrix}\begin{bmatrix}1&0&0&0&0&0&\Delta_{17}&\cdots&\Delta_{1n}\\ 0&1&0&0&0&0&0&\cdots&0\\ 0&0&1&0&0&0&0&\cdots&0\\ 0&0&0&1&0&0&0&\cdots&0\\ 0&0&0&0&1&0&\Delta_{57}&\cdots&\Delta_{5n}\\ 0&0&0&0&0&1&\Delta_{67}&\cdots&\Delta_{6n}\\ \end{bmatrix}$
@@
The columns needed to span $\mathbf{F}^{n,m}\circ\mathbf{S}^{n,d}$ are the same as those needed for $\mathbf{F}^{n,m}\circ\mathbf{F}^{n,m}$, with the same sparsity pattern. Therefore, we will have
$\mathrm{rank}(\mathbf{F}^{n,m}\circ(\mathbf{S}^{n,d}+\mathbf{F}^{n,m}))=\mathrm{rank}(\mathbf{F}^{n,m}\circ\mathbf{S}^{n,d})\leq\min(2m,n)$ (18)
@@
$=-\frac{1}{2}J^{n}(\mathbf{D}^{n,d,0}+\mathbf{F}^{n,m}\circ(\mathbf{S}^{n,d}+\mathbf{F}^{n,m}))J^{n}$
$=-\frac{1}{2}J^{n}(\mathbf{1}\,\mathrm{diag}(\mathbf{X}^{\top}\mathbf{X})^{\top}-2\mathbf{X}^{\top}\mathbf{X}+\mathrm{diag}(\mathbf{X}^{\top}\mathbf{X})\mathbf{1}^{\top})J^{n}-\frac{1}{2}J^{n}(\mathbf{F}^{n,m}\circ(\mathbf{S}^{n,d}+\mathbf{F}^{n,m}))J^{n}$
@@
The remaining term is $\frac{1}{2}J^{n}(\mathbf{F}^{n,m}\circ(\mathbf{S}^{n,d}+\mathbf{F}^{n,m}))J^{n}$. By rank inequality,
$\mathrm{rank}(J^{n}(\mathbf{F}^{n,m}\circ(\mathbf{S}^{n,d}+\mathbf{F}^{n,m}))J^{n})\leq\min(\mathrm{rank}(\mathbf{F}^{n,m}\circ(\mathbf{S}^{n,d}+\mathbf{F}^{n,m})),\mathrm{rank}(J^{n}))=\min(2m,n-1)$ (24)
@@
Appendix C Proof of Proposition 3.3
@@
In terms of rank, the noise acts as many small faults on each satellite, with the same sparsity structure as $\mathbf{F}^{n,m}$ in Proposition 3.1, almost surely with $m=n$ since there is zero probability mass that the randomly sampled noise is exactly zero or exactly cancel out. Using Proposition 3.2,